# Logarithmic Finite-Size Scaling of the Four-Dimensional Ising Model

### v1

@@ III.2 Thermal scaling behaviors
We next explore the FSS of the size of the largest and second-largest loop clusters $F_{1},F_{2}$ in the loop Ising model, in particular, to examine the form of logarithmic corrections. Different from the FK clusters, size of loop clusters are thermal quantities, with the fractal dimension equal to the thermal exponent $y_{t}$, as numerically observed for $d>4$ in Ref. xiao2023finitesize. Thus at $d=4$, we expect $F_{1},F_{2}\sim L^{2}$, accompanied by multiplicative logarithmic corrections. To find the form of logarithmic corrections, again we plot the rescaled largest loop cluster size $\tilde{F}_{1}:=F_{1}/L^{2}$ versus $\ln{L/L_{0}}$ in Fig. 3(b1) with $L_{0}$ fixed at $0.59$. Except for the small systems which suffer strong finite-size effect, data with large system collapse nicely on the dashed line with slope $1/6$. Thus, asymptotically, our data support the scaling $F_{1}\sim L^{2}(\ln{L/L_{0}})^{1/6}$. To further confirm, we also plot in Fig. 3(b2) the $\tilde{F}_{1}^{6}$ versus $\ln L$; in this way, the uncertainty from fixing the constant $L_{0}$ is removed. The clear linear relation demonstrates that $F_{1}\sim L^{2}(a_{1}\ln L+a_{2})^{1/6}$. Thus, our data strongly suggest that $F_{1}\sim L^{y_{t}}(\ln L)^{\hat{y}_{t}}$ with $\hat{y}_{t}=1/6$. We note that, in comparison with the specific heat $c_{\textsc{e}}$, the background term $a_{0}$ in $F_{1}\sim a_{0}+a_{1}L^{2}[\ln(L/L_{0})]^{1/6}$ plays a less important role.
@@ IV Conclusion
In this work, we carry out a systematic study on the logarithmic corrections in the finite-size scaling (FSS) of the four-dimensional (4D) Ising model under two geometric representations, i.e., the Fortuin-Kasteleyn (FK) random-cluster and the loop representations. The study of the 4D Ising model is significant due to its connection to the O$(n)$ model, the scalar sector of the standard model, and its relevance in various condensed matter systems, where logarithmic correction behaviors are expected to manifest at the three-dimensional quantum critical point (QCP) zinn2021quantum; Higgs1964; higgsmodeinCMP.
@@
In both the FK Ising and the loop Ising models, our data show that the FSS of the critical susceptibility is consistent with the expected scaling $\chi\sim L^{2y_{h}-d}(\ln L)^{2\hat{y}_{h}}$ with $y_{h}=3$ and $\hat{y}_{h}=1/4$. Moreover, in the FK Ising model, our data suggest that the size of the largest FK cluster scales as $C_{1}\sim L^{y_{h}}(\ln L)^{\hat{y}_{h}}$, and the size of the second-largest cluster scales as $C_{2}\sim L^{y_{h}}(\ln L)^{\hat{y}_{h2}}$ with $\hat{y}_{h2}=-1/4$ a new correction exponent which has no theoretical prediction. In the loop Ising model, the specific heat is observed to scale as $c_{\textsc{e}}\sim a_{0}+a_{1}\left[\ln(L/L_{0})\right]^{2\hat{y_{t}}}$ with $\hat{y_{t}}=1/6$, which numerically confirm the logarithmic divergence of the critical specific heat of the 4D Ising model. For the size of the largest loop cluster, our data suggest $F_{1}\sim L^{y_{t}}(\ln L)^{\hat{y}_{t}}$ with $\hat{y}_{t}=1/6$.
@@
We finally provide a simple explanation for why the logarithmic divergence of the specific heat is much easier to be numerically observed in the loop Ising model. From the conjectured FSS ansatz (10), the predicted FSS for the specific heat can be written as in Eq. (15), in which the constant $a_{0}$ is the background term and the $a_{1}$-term corresponds to the modified CG term. If $a_{0}$ is comparable to $a_{1}$ and $L_{0}$ is unknown, then it is hard to numerically extract the value of the exponent $\hat{y_{t}}$ from the data. But in the loop Ising model, our data show that the background term in $c_{\textsc{b}}$ is very small, so the logarithmic correction can be easily extracted using Fig. 3(a2). Once the logarithmic correction is determined for $c_{\textsc{b}}$, the scaling of the specific heat follows directly from Eq. (14). More convincing evidence can be seen from the size of the largest loop cluster $F_{1}$. As proposed in Ref. xiao2023finitesize, the FSS of $F_{1}$ completely follows the CG asymptotics for $d>d_{c}$. If this picture also holds at $d=d_{c}$, then at $d_{c}$ it follows from Eq. (10) that $F_{1}\sim a_{0}+a_{1}L^{y_{t}}\left[\ln(L/L_{0})\right]^{\hat{y}_{t}}$, without the effect from the GFP term. Thus, if we study the ratio $F_{1}/L^{y_{t}}$, then in comparison with the logarithmic term, the background $a_{0}$-term becomes subdominant with order $L^{-y_{t}}$. In $c_{\textsc{b}}$, the effect of the background to the logarithmic term is also subdominant but of order $(\ln L)^{-\hat{y}_{t}}$, much larger than $L^{-y_{t}}$. In other words, in $F_{1}$, the background term has much weaker effect to the estimate of the exponent $\hat{y}_{t}$, and thus the result from $F_{1}$ is more convincing.

### v2

@@ III.2 Thermal scaling behaviors
We next explore the FSS of the size of the largest and second-largest loop clusters $F_{1},F_{2}$ in the loop Ising model, in particular, to examine the form of logarithmic corrections. Different from the FK clusters, size of loop clusters are thermal quantities, with the fractal dimension equal to the thermal exponent $y_{t}$, as numerically observed for $d>4$ in Ref. Xiao et al. (2024). Thus at $d=4$, we expect $F_{1},F_{2}\sim L^{2}$, accompanied by multiplicative logarithmic corrections. To find the form of logarithmic corrections, again we plot the rescaled largest loop cluster size $\tilde{F}_{1}:=F_{1}/L^{2}$ versus $\ln{L/L_{0}}$ in Fig. 3(b1) with $L_{0}$ fixed at $0.59$. Except for the small systems which suffer strong finite-size effect, data with large system collapse nicely on the dashed line with slope $1/6$. Thus, asymptotically, our data support the scaling $F_{1}\sim L^{2}(\ln{L/L_{0}})^{1/6}$. To further confirm, we also plot in Fig. 3(b2) the $\tilde{F}_{1}^{6}$ versus $\ln L$; in this way, the uncertainty from fixing the constant $L_{0}$ is removed. The clear linear relation demonstrates that $F_{1}\sim L^{2}(a_{1}\ln L+a_{2})^{1/6}$. Thus, our data strongly suggest that $F_{1}\sim L^{y_{t}}(\ln L)^{\hat{y}_{t}}$ with $\hat{y}_{t}=1/6$. We note that, in comparison with the specific heat $c_{E}$, the background term $a_{0}$ in $F_{1}\sim a_{0}+a_{1}L^{2}[\ln(L/L_{0})]^{1/6}$ plays a less important role.
@@ IV Conclusion
In this work, we carry out a systematic study on the logarithmic corrections in the finite-size scaling (FSS) of the four-dimensional (4D) Ising model under two geometric representations, i.e., the Fortuin-Kasteleyn (FK) random-cluster and the loop representations. The study of the 4D Ising model is significant due to its connection to the O$(n)$ model, the scalar sector of the standard model, and its relevance in various condensed matter systems, where logarithmic correction behaviors are expected to manifest at the three-dimensional quantum critical point (QCP) Zinn-Justin (2021); Higgs (1964); Pekker and Varma (2015).
@@
In both the FK Ising and loop Ising models, our data indicate that the finite-size scaling (FSS) of the critical susceptibility aligns with the expected scaling $\chi\sim L^{2y_{h}-d}(\ln L)^{2\hat{y}_{h}}$, where $y_{h}=3$ and $\hat{y}_{h}=1/4$. This excellent consistency validates the effectiveness of our numerical simulation, suggesting that the deviation between the coupling we simulated at and the genuine critical coupling has a negligible impact on our results. Moreover, in the FK Ising model, our data suggest that the size of the largest FK cluster scales as $C_{1}\sim L^{y_{h}}(\ln L)^{\hat{y}_{h}}$, and the size of the second-largest cluster scales as $C_{2}\sim L^{y_{h}}(\ln L)^{\hat{y}_{h2}}$ with $\hat{y}_{h2}=-1/4$ a new correction exponent which has no theoretical prediction. In the loop Ising model, the specific heat is observed to scale as $c_{\textsc{e}}\sim a_{0}+a_{1}\left[\ln(L/L_{0})\right]^{2\hat{y_{t}}}$ with $\hat{y_{t}}=1/6$, which numerically confirm the logarithmic divergence of the critical specific heat of the 4D Ising model. For the size of the largest loop cluster, our data suggest $F_{1}\sim L^{y_{t}}(\ln L)^{\hat{y}_{t}}$ with $\hat{y}_{t}=1/6$.
@@
We finally provide a simple explanation for why the logarithmic divergence of the specific heat is much easier to be numerically observed in the loop Ising model. From the conjectured FSS ansatz (10), the predicted FSS for the specific heat can be written as in Eq. (15), in which the constant $a_{0}$ is the background term and the $a_{1}$-term corresponds to the modified CG term. If $a_{0}$ is comparable to $a_{1}$ and $L_{0}$ is unknown, then it is hard to numerically extract the value of the exponent $\hat{y_{t}}$ from the data. But in the loop Ising model, our data show that the background term in $c_{\textsc{b}}$ is very small, so the logarithmic correction can be easily extracted using Fig. 3(a2). Once the logarithmic correction is determined for $c_{\textsc{b}}$, the scaling of the specific heat follows directly from Eq. (14). More convincing evidence can be seen from the size of the largest loop cluster $F_{1}$. As proposed in Ref. Xiao et al. (2024), the FSS of $F_{1}$ completely follows the CG asymptotics for $d>d_{c}$. If this picture also holds at $d=d_{c}$, then at $d_{c}$ it follows from Eq. (10) that $F_{1}\sim a_{0}+a_{1}L^{y_{t}}\left[\ln(L/L_{0})\right]^{\hat{y}_{t}}$, without the effect from the GFP term. Thus, if we study the ratio $F_{1}/L^{y_{t}}$, then in comparison with the logarithmic term, the background $a_{0}$-term becomes subdominant with order $L^{-y_{t}}$. In $c_{\textsc{b}}$, the effect of the background to the logarithmic term is also subdominant but of order $(\ln L)^{-\hat{y}_{t}}$, much larger than $L^{-y_{t}}$. In other words, in $F_{1}$, the background term has much weaker effect to the estimate of the exponent $\hat{y}_{t}$, and thus the result from $F_{1}$ is more convincing.